### v2

# Linguistic Generalizations are not Rules: Impacts on Evaluation of LMs

Abstract
Linguistic evaluations of how well LMs generalize to produce or understand novel text often implicitly take for granted that natural languages are generated by symbolic rules. Grammaticality is thought to be determined by whether sentences obey such rules. Interpretation is believed to be compositionally generated by syntactic rules operating on meaningful words. Semantic parsing is intended to map sentences into formal logic. Failures of LMs to obey strict rules have been taken to reveal that LMs do not produce or understand language like humans. Here we suggest that LMs’ failures to obey symbolic rules may be a feature rather than a bug, because natural languages are not based on rules. New utterances are produced and understood by a combination of flexible, interrelated, and context-dependent constructions. We encourage researchers to reimagine appropriate benchmarks and analyses that acknowledge the rich, flexible generalizations that comprise natural languages.
Linguistic Generalizations are not Rules: Impacts on Evaluation of LMs
Leonie Weissweiler    Kyle Mahowald The University of Texas at Austin {weissweiler,kyle}@utexas.edu                           Adele E. Goldberg    Princeton University    adele@princeton.edu
1 Introduction
Constructions: Learned Pairings of Form and Function Rules
(Partially-filled) words, common and rare (partially-filled) schemata Common abstract patterns
Wide range of functions Only abstract functions
Combination of open slots and/or fixed lexical units Include only variables
Context-sensitive (and plentiful) Context-free (and few)
Inter-related within a complex network Unstructured list
Sensitive to similarity and frequency Insensitive to similarity or frequency
Slots constrained in open-ended range of ways Variables constrained by gram. category
Table 1: Differences between constructions and rules
How well do large Language Models (LMs) generalize beyond their training data? Much work on this question has presumed that generalizations require symbolic rules for syntax and semantics that generate acceptable new forms and compositional meanings. Rules are invoked to explain that if you learn a new modifier (‘blonky’) and a new count noun (‘gravimin’), a compositional rule could predict that ‘a blonky gravimin’ is a gravimin that is blonky. In what follows, we use “rule” to refer to context-free generalizations that contain variables, to be instantiated by any instance of a general type. In the domain of language, rules are assumed to be uninfluenced by frequency, similarity, or context Pinker (1999). Rules of this type are relied on in math, logic, and programming languages.
Because earlier statistical models (e.g., n-gram or Markov models) seemed unable to generalize fully or capture non-local dependencies Chomsky (1957), rules seemed to many to be the only game in town for human language. After all, if a standard bigram model hadn’t seen ‘blonky gravimin’ before, it would be unable to interpret it. Influential thinkers argued that neural networks, which did not involve rules, would never be appropriate models of human cognition for this reason Fodor and Pylyshyn (1988); Pinker and Prince (1988); Marcus (1998); Fodor and Lepore (2002); Marcus (2001); Calvo and Symons (2014). Yet GPT-4o, for example, not only recognizes ‘a blonky gravimin’ is a noun phrase, it explicitly guesses several naturalistic interpretations, e.g., ‘A person or act that awkwardly and absurdly pretends to be serious.’
Critically, today’s LMs arose from statistical, distributional parallel models Mikolov et al. (2013); Rumelhart et al. (1986) rather than rule-based natural language technologies. They do not rely on hard-coded rules, yet their ability to produce coherent, naturalistic language and respond appropriately is unparalleled by purely symbolic systems Piantadosi (2024); Goldberg (2024); Weissweiler et al. (2023); Hofmann et al. (2025).
Nonetheless, an assumption that generalizations are equivalent to rules continues to motivate many evaluations of syntax, meaning, and their compositional combination: e.g., Natural Language Inference Bowman et al. (2015), Semantic Parsing Palmer et al. (2005); Reddy et al. (2017), tests of binary grammatical acceptability Warstadt et al. (2019); Dentella et al. (2023) and rule-based compositionality Kim and Linzen (2020). Together, such tasks made up more than half of the GLUE benchmark Wang et al. (2018), created to evaluate language models on their skill at being “general, flexible, and robust.” Lackluster performance on rule-based tasks in the early days of LMs was taken to imply that the models did not use language the way people do and were instead merely imitating shallow surface patterns Lake and Baroni (2018); Kim and Linzen (2020); Weißenhorn et al. (2022); Bolhuis et al. (2023). In a survey of 79 NLP researchers, McCurdy et al. (2024) reported that 87% believed LMs were not sufficiently compositional and a sizable proportion (39%) believed explicit discrete symbolic rules were required.
Evaluations of early LMs’ challenges using algebraic or logical rules did expose certain shortcomings in their ability to reason abstractly and solve math problems (e.g., Mahowald et al., 2024). At the same time, LM’s ability to produce and respond to natural languages naturalistically is hard to overstate (e.g., Coil and Shwartz, 2023). We therefore suggest that rule-based evaluations have been over-emphasized in the domain of natural language production and comprehension. More nuanced metrics, measuring gradient judgments and context-dependent interpretations are more valuable (e.g., Juzek, 2024; Hu et al., 2024a).
Rules are not sufficient for generalization. For LMs to use language like humans, they require interpretations that are far richer than rules can provide for thousands of collocations, conventional metaphors, idioms, and context-dependent interpretations. Abstract grammatical patterns likewise regularly convey semantic and/or pragmatic information that restrict their contexts of use and interpretation.
Just as LMs appear to do Tseng et al. (2022), we propose that each natural language requires mastering a network of hundreds of thousands of context-dependent, gradient, flexible schemata (constructions, see §5), which often contain ‘slots’ that constrain their fillers and how those fillers are interpreted. Constrained slots allow for new combinations, flexibly adapted in context. For instance, the phrase ’<time period> ago’ can coerce a temporal interpretation of filler phrases that do not designate time periods (e.g., ‘three rest stops ago’). Rather than rule-based composition of meaning, composition-by-construction allows constructions to contribute meaningfully to interpretation in ways that range from quite specific to largely abstract.
A list of differences between rules and constructions is provided in Table 1, and we explain each row in §5. Constructions anticipate that different languages or dialects will provide speakers with different ConstructionNets, so that cross-linguistic differences can be captured naturally.
We feel that context-free, fully abstract rules are not necessary, as we suggest in §5 that constructions do what rules are meant to do and more. At the same time, the approach we advocate need not eschew the possibility of context-free, fully abstract rules that apply within a particular language, or even, in the limiting case, across languages. If such cases exist, they can be treated as a fully abstract construction Jackendoff (2002). Note that Jackendoff (2002) proposes a rule Verb + Particle to capture the syntax of English complex verbs, but no one would argue that the meanings of complex verbs arise from a compositional rule (e.g., one can look up a number or look down on someone but not ?look up on someone or ?look down a number). We are simply advocating for an increased focus on the extent to which and how LMs manage to produce and comprehend human-like natural languages in all their context-specificity and complexity.
Many of our theoretical points are not new. While early AI relied on algebraic rules Minsky and Papert (1969); Lenat (1995), many researchers soon realized that rules were too brittle to scale up beyond highly restricted domains such as artificial block worlds Winograd (1980). Neural network researchers have continuously argued in favor of a single representational system and against the usefulness of rules, primarily in the domain of words and inflectional morphology (e.g., Rumelhart et al., 1986; Rogers and McClelland, 2004; Elman, 2009; Christiansen and Chater, 1999; MacDonald et al., 1994; McClelland, 2015).
Our contribution is to review leading paradigms used in LM evaluations of syntax (§2), semantics (§3), and compositionality (§4). We explain how rules are implicitly assumed in many impactful evaluations, briefly describing how the assumptions arose, and why they are problematic. We propose an alternative theoretical basis (§5), encouraging the field to evaluate LMs on the extent to which they learn and represent the complex network of constructions that comprises each language and how constructions are combined and generalized to produce new messages in new contexts.
2 Syntactic Rules in LM Evaluations
An assumption that natural languages are generated by rules such as phrase structure rules, movement rules, or an operation ‘Merge’ has been assumed by most versions of generative grammar since Chomsky (1957). Syntactic rules are intended to operate on broad and clearly defined grammatical categories (e.g., nouns, verbs, adjectives) or relations (e.g., subject, object), which are in turn combined in rule-like fashion to create larger units (e.g., noun phrases, verb phrases, adjective phrases). Rules are claimed to be uninfluenced by frequency, similarity, meaning, or context Pinker (1999).
Grammaticality judgment tasks
Human judgments on sentences are gradient rather than binary, and demonstrably depend on frequency, plausibility, complexity, memory demands, potential alternatives, and context Grodner and Gibson (2005); Schütze and Sprouse (2013); Robenalt and Goldberg (2015); Gibson and Hickok (1993); Fang et al. (2023). The amount of exposure to written language or linguistic theory also influences judgments. For instance, Dąbrowska (2010) found that laypeople’s judgments on sentences containing long-distance dependencies were more sensitive to lexical content than linguists’ judgments were.
Nonetheless, LMs’ language skills are often evaluated on binary grammaticality judgments on sentences Dentella et al. (2024, 2023); Warstadt et al. (2019) Even sentences included in linguistic textbooks, which one might presume to have clear-cut judgments, in reality are judged gradiently by people Juzek (2024).
The fact that human judgments are gradient can have profound consequences on evaluations. For instance, Dentella et al. (2023) compared humans and LMs against predetermined binary acceptability labels, reporting that LMs’ performance correlated poorly. However, comparing gradient perplexity measures with the same human judgments revealed a strong positive correlation Hu et al. (2024b). Using perplexity measures for models (as well as allowing humans to provide ordinal or gradient judgments) is a step in the right direction Hu et al. (2024b); Juzek (2024).
Dependency Parsing as Evaluation
Parsing text for universal dependencies (UD, de Marneffe et al., 2021) has become a well-established task for evaluating models Zeman et al. (2017, 2018), and since Hewitt and Manning (2019) showed BERT Devlin et al. (2019) to be somewhat skilled in UD, it has became the default operationalization of syntax in the NLP world Amini et al. (2023); Kryvosheieva and Levy (2025); Müller-Eberstein et al. (2022) and in discussions of inductive biases Lindemann et al. (2024); Glavaš and Vulić (2021). UD annotations are partially determined by semantics, and they are based on lexical items, which makes them closer to the approach advocated here rather than abstract phrase structure rules. However, UD analyses presume a universal set of grammatical relations, which is problematic, because not all languages employ the same constructs. That is, there is no universally valid way to define or test for nouns, verbs, subjects, or direct objects (e.g., Croft, 2001). Moreover, UD requires an asymmetric relationship between a ’head’ and dependent, yet the long tail of language includes headless constructions (e.g., the Xer, the Yer construction) Michaelis (2003) and co-headed constructions (e.g., phrasal verbs, conjunctions, and idioms). Therefore, evaluating LMs on UD annotations may set a standard that is misaligned wth aspects of natural languages. Most importantly, LMs produce and interpret meaningful language to varying degrees. Yet, since UD is a syntactic annotation schema, it provides no means of generating meaningful language nor assigning meaningful interpretations.
3 Semantic Rules in LM Evaluation
Formal logic was developed as a branch of mathematics, used to prove mathematical and philosophical theorems and identify provability gaps Frege (1918); Russell (1905); Gödel (1931). It was based on algebraic rules operating on clear-cut and broadly defined categories. Notably, logicians did not generally assume nor endorse using formal logic to represent the meanings of natural language utterances Carnap (1937); Baker and Hacker (1986), recognizing that natural languages differ from formal logic in many ways. For instance, logic treats and and but as equivalent. It does not provide a natural way to capture commands or questions (Austin, 1975), nor does it naturally distinguish presuppositions from assertions Strawson (1967). Finally, formal logic is not intended to capture effects of context Wittgenstein (1953); Russin et al. (2024).
Natural Language Inference
Natural Language Inference tasks label the second of two sentences as an entailment, contradiction, or neutral. NLI tasks were originally used to train models Wang et al. (2019); Dagan et al. (2006); Nie et al. (2020). Today, they are used as zero-shot evaluation metrics to assess natural language understanding Zhou et al. (2024); McCoy et al. (2019). Entailment and contradiction are key concepts in mathematical proofs. However, the centrality of entailment and contradiction in language understanding can be overstated (e.g., “The semantic concepts of entailment and contradiction are central to all aspects of natural language meaning. ” Bowman et al. (2015) ) (emphasis added, see also Katz, 1972; van Benthem, 2008). Necessary and plausible inferences are a critical aspect of natural language understanding, but they are highly dependent on context, as well as the interlocutors’ general communicative goals.
Humans’ goal is to make sense of others’ messages, so we assume others are trying to be relevant and helpful and do our best to assign coherent meanings to all utterances Grice (1975). For example, outside of logic classes or heated arguments, people rarely conclude that two statements made by the same person are contradictory. If someone utters: ‘The boy is depressed but he is not DEPRESSED’, listeners do not throw up their hands and shout ‘contradiction!’. Instead, they may infer that the boy in question is only somewhat, and not extremely, depressed, or ask to learn more. Humans also assign distinct context-dependent interpretations to apparent tautologies such as ‘Either it’s alive or it’s not’ and ‘If it snows, it snows.’ NLI tasks that rely on judging contradictions or entailments may over- or under-estimate how well LMs’ understanding of natural language aligns with humans’, particularly when binary judgments are required (cf. Dentella et al., 2024).
We need evaluation metrics to take context and communicative goals into account and to allow to allow gradient judgments. An example of the type of evaluation we endorse can be found in the underappreciated NOPE testbed created by Parrish et al. (2021). The authors selected 10 distinct constructions that trigger presuppositions, and curated 100+ instances of each one, based on naturally occurring examples. Each stimulus included two preceding sentences for context. The authors then collected gradient judgments from human raters that allowed them to use their ‘background knowledge about how the world works’ and compared the accuracy of several models, with several controls in place, which strikes us as the right approach.
Semantic Parsing
Banarescu et al. (2013) introduced abstract graphical meaning representations (AMR) for sentential meaning that importantly includes aspects of lexical semantics. It was created to offer a repository of structured meanings to be used for evaluating understanding in LMs, a worthwhile goal (Li et al., 2023; Qiu et al., 2022; Shaw et al., 2021, see also §4). At the same time, the representation explicitly ignores ’syntactic idiosyncracies.’ For example, ‘he described her as a genius’ and ‘his description of her: genius’ are assigned the same AMR. Yet the former is unambiguously about her intellect, while the latter may instead be used to compliment his cleverly tactful description. More generally, simplifications of distinctions made in a natural language commonly result in a loss in meaning, since two utterances are rarely fully interchangeable. Languages rarely offer two ways to express the identical message in context.
4 Compositionality
As computer coding languages became more and more widespread, rule-based syntax and semantics took root in linguistics. A Principle of Compositionality states that the semantics of a sentence is determined by the meanings of the words and the syntactic rules used to combine them Montague (1970); Partee (1984); Dowty (1979); Jackendoff (1992); Fodor and Lepore (2002). It is intended to be a bottom-up process: syntactic rules combine words, which have determinant meanings. Fodor and Pylyshyn (1988) make this clear: “a lexical item must make approximately the same semantic contribution to each expression in which it occurs”. That is, context may not influence the interpretation of words in a top-down manner; therefore downstream inferences are required to address the fact that interpretations do depend on context. Realizing this, like Carnap and Frege before him, Fodor and Pylyshyn (1988) acknowledge: “It’s uncertain exactly how compositional natural languages actually are.”
Nonetheless, compositionality is often taken as a truism, based on the standard argument for it summarized below.
People tend to agree on the interpretation of new sentences. $\Rightarrow$ There must be some set of rules that determine the meaning of new sentences.
One can agree with the premise without accepting its consequent.
In particular, people generally agree on the interpretations of pointing gestures and novel words as well as sentences, and yet shared interpretations must be gleaned from non-linguistic context in the case of pointing gestures, and from a combination of linguistic and non-linguistic context in the case of novel words. Shared interpretation of sentences likewise comes in part from linguistic and non-linguistic context.
Consider the sentence, ‘the Persian cat is on the mat.’ If the speaker’s goal is simply to help someone find the furball, there need be no commitment to the cat being a thoroughbred Persian breed nor to the cat being wholly on, rather than adjacent to, the mat. Or, comprehenders may appreciate the statement is ironic if the cat is hairless.
Cases that might seem amenable to a rule often turn out to require a good deal of item-specific memory. For instance, a compositional rule involving set-intersection may seem appealing for ‘<color term> noun’ combinations in the domain of artificial block worlds (e.g., a green cube is something that is both a cube and green). However, violations of such rules abound: green tea is more yellow than green, and Cambridge blue is actually green. Even more common are instances that evoke richer meanings than predicted by any algebraic rule: e.g., a green light implies that forward motion or progress is permitted, and a green card provides a path toward citizenship in the US (or ought to). The meanings of familiar collocations and idioms are typically not fully determined by general rule.
Another issue is that rules massively over-genenerate, which is why they are not used for production. That is, rules predict all manner of odd locutions Pawley and Syder (1983); Sag et al. (2002): e.g., ‘Meeting you is pleasing to me’; ‘The tall winds hit the afraid boy’; ‘Explain them the problem.’ Humans are sensitive to the frequencies of various types of word combinations and judge formulations unnatural whenever there exists a more conventional way to express the intended message in context (e.g., Goldberg, 2019).
Evaluating LMs for Compositionality
Compositionality benchmarks combine elements from syntactic and semantic evaluations. Kim and Linzen (2020)’s compositional generalization challenge (COGS) tested whether models could translate any sentence generated by a small set of syntactic rules into formal semantics. For instance, trained on representations of ‘the girl,’ ‘the cat,’ ‘the hedgehog,’ ‘the cat loves the girl,’ and ‘the hedgehog sees the cat,’ and so on, the model was tested on how well it predicted a formal semantic representation for ‘The girl loves the hedgehog.’ However, note that if ‘mosquitoes’ is substituted for ‘the cat,’ different interpretations of ‘love’ are evoked (‘Mosquitoes love the girl’ vs. ‘The girl loves mosquitoes’), not to mention markedly different degrees of plausibility. The authors also anticipated generalizations from sentences like ‘Jane gave the cake to John’ to ‘Jane gave John the cake,’ and the models were found to perform poorly. Yet the two sentences differ in terms of information structure Bresnan and Ford (2010), and the relative frequencies and similarities of verbs witnessed in each version Leong and Linzen (2024); Ambridge et al. (2014); Hawkins et al. (2020). Insofar as natural language is not amenable to logical representations, failures of LMs to map to such representations may be consistent with human interpretation.
Other compositionality benchmarks adopt NLI tasks, which commonly presume interpretation is determined by rules. For example, in the context of robotic agents interpreting instructions, Lake and Baroni (2018, p.1) state:
Humans can understand and produce new utterances effortlessly, thanks to their compositional skills. Once a person learns the meaning of a new verb ‘dax’, he or she can immediately understand the meaning of ‘dax twice’…
The robotic agents struggled to interpret the rule-based command, though it was appropriate in the narrow domain tested. Notably, the rule does not extend to a broader swath of language. For instance, unbounded actions are not countable, so if ‘twice’ appears at all, it is likely followed by a comparative phrase (e.g., ‘work twice as hard’) and a very different meaning than performing an action two times. Other cases require knowledge of specific combinations: ‘to think twice’ means ’to hesitate’ and ‘going twice’ tends to evoke the context of an auction. Familiar phrases with meanings not fully captured by compositional rules are common: By one estimate, we learn tens of thousands of them Jackendoff (2002). Importantly, we largely agree on their interpretations, even though each means something more or different than predicted simply by the words and their syntactic combination. In this way, phrasal combinations regularly involve subregularities or item-specific interpretations not predicted by a general algebraic rule.
Another example comes from the seemingly innocuous algebraic rule: “If X is more Y than Z, then Z is less Y than Z, irrespective of the specific meanings of X, Y, and Z” (Dasgupta et al., 2020, p.5). This is meant to capture that ‘Anne is more cheerful than Bob’ should both contradict ‘Anne is less cheerful than Bob’, and entail ‘Bob is less cheerful than Anne.’ NLI models that failed to draw these inferences were considered lacking. Yet natural language rarely relies on free variables. The content of X, Y, and Z matters. No one would infer that because Annex is more cheerfuly than carefulz, that ‘Carefulz is less cheerfuly than Annex.’ Perhaps more importantly, if a speaker uttered ‘Anne is higher than Bob and Bob is higher than Anne,’ listeners would likely infer either that Bob climbed above Anne in the time it took to utter the first clause or that Bob has been smoking. We have so far argued against symbolic abstract rules for natural languages. Below we suggest a positive alternative.
5 Constructions
This section briefly explains an alternative to positing a set of abstract symbolic rules, namely a network of interrelated constructions. Constructions differ from rules in each dimension cited in Table 1 and briefly explained below. For more comprehensive introductions, see Hoffmann and Trousdale (2013) and Hoffmann (2022).
(Partially-filled) Words, Common & Rare Schemata
A ‘construction’ is any learned association between a formal pattern and a range of related functions. This simple definition treats words, idioms, rare and common grammatical patterns as constructions. Formal attributes may include phonology, lexical content, semantic or formal categories, word order, discontinuous elements, and/or intonation. Any construction may include one or more constrained open ‘slots’ (Table 2).
Wide Range of Functions
Constructions’ functions vary widely: Constructions convey rich, specific, contentful meaning in the case of words, collocations, and idioms. A plethora of other constructions are productive but constrained in a wide variety of semi-specific ways; argument structure constructions convey ‘who did what to whom’; discourse structuring constructions indicate which parts of an utterance are at-issue or backgrounded. A range of constructions exist to ask questions, express surprise or disapproval, greetings or gossip. Any level of construction can be associated with specific registers, genres, and/or dialects.
Construction Type Examples
Words pregame; running; nevertheless; ago
[…]
Collocations, fixed idioms high winds; jump the shark
[…]
Partially-filled argument structure constructions give <recipient> a call
Argument structure constructions Intransitive; Caused-motion; Double Object; Resultative
Discourse-structuring constructions get-passive; information questions; it-clefts; relative clauses
Table 2: Example constructions at varying levels of complexity and abstraction
Sensitive to Similarity and Frequency
Language users are sensitive to the frequencies of constructions. For instance, the passive construction is used far more frequently in Turkish than English and young Turkish speakers learn the construction far earlier than English-speaking children Slobin (1986). Constructions are also influenced by similarity: Instances of a construction prime instances of the same or closely related construction (e.g., Du Bois, 2014; Pickering and Ferreira, 2008).
Productive Constructions May Include Fixed Lexical Units
As shown in Table 2, syntax, semantics, and morphology are interrelated rather than assigned distinct levels. This is useful because even productive hierarchical constructions often include particular words and semantic constraints. For example, an English construction that implies real or metaphorical motion allows a wide range of verbs but requires the particular noun ‘way’ (‘He charmed his way into the meeting.’).
Plentiful and Context-Sensitive
The broad definition of constructions as pairings of form and function, including words and grammatical patterns, is another difference between constructions and rules. Constructions are far richer and more plentiful than the class of rules is commonly envisioned to be, as they combine aspects of form and function. Constructions also do far more work than rules since they capture frequency information and contextual constraints, while rules are presumed to be context-free and uninfluenced by frequency (e.g., Pinker, 1999).
Interrelated System, Not Unstructured List
Unlike rules, which are commonly presented as unstructured lists, constructions comprise a network of interrelated statistical patterns. This allows for the fact that languages have families of related constructions. It also allows for the simple fact that productive constructions simultaneously co-exist with specific conventional instances. For instance, the English ‘double object’ construction is productive, and speakers are also familiar with dozens of conventional instances (e.g., ‘give <someone> the time of day’, ‘throw <someone> a bone’).
Construction Slots Are Constrained
The open ‘slots’ of constructions are constrained in a wide variety of ways. For instance, the English double-object construction can appear with a wide range of verbs, but prefers simple verbs to those that sound Latinate (e.g., ‘She told them something’ vs. ¿She proclaimed them something’). The English comparative suffix ‘-er’ (e.g., ‘calmer’, ‘quicker’) is available for most single-syllable adjectives that allow a gradient interpretation, but it is not used with past participle adjectives (? ‘benter’).
An Example
Consider ‘X is the new Y’. It is productive and can be used to create new utterances, e.g., ‘Semiconductor chips are the new oil.’ As is typical of productive constructions, the generalization co-exists with several familiar instances (e.g., ‘50 is the new 40’; ‘Orange is the new black’). The construction is not an algebraic rule: Its slots, indicated by X and Y, are not variables that range freely over fixed syntactic categories. Instead, ‘X’ must be construed (playfully) as currently functioning in the culture as ‘Y’ used to. Therefore, not all combinations of slot fillers make sense: (e.g., ? ‘Orange is the new oil’). Adding a parallelism constraint between X and Y is insufficient since ‘103 is the new 101’ would also require an unusual context to make sense. Finally, instances of the construction are not amenable to a general compositional rule, nor can it be translated into formal logic, which would presumably treat ‘Orange is the new black’ as equivalent to ‘Black is the old orange,’ which does not conventionally evoke the same meaning.
6 Implications Beyond Natural Language
Outside of natural language, even in domains that are rule-like by design, certain types of non-rule-based interpretations exist, likely due to the interface from natural language. For instance, LMs have been found unreliable at drawing the following inference, which the authors dubbed the reversal curse: “if ‘A is B’ […] is true, then ‘B is A’ follows by the symmetry property of the identity relation" (Berglund et al., 2023, p. 2).
Why are LMs prone to the reversal curse? Although the quote above is stated in natural language, it does not apply to natural language sentences, which are rarely reversible with the same interpretation: e.g., ‘A mental illness is the same as a physical illness’ means something very different than ‘A physical illness is the same as a mental illness’ (see also Tversky, 1977; Talmy, 1975). Even simple conjunctions are not generally reversible in natural language. For instance, ‘night & day’ and ‘day & night’ are both acceptable, but their interpretations differ: the former conveys a stark contrast (e.g., ‘as different as night and day’), the latter suggests a relentless activity or process (e.g., ‘he worried day and night’). In summary, it is perhaps reasonable to expect truly symmetric knowledge to be reversible. But LMs are trained on natural language, which is not symmetric.
7 New Directions for Evaluation
Natural languages involve complex and context-sensitive systems of constructions, which vary from being wholly fixed to highly abstract and productive (recall Table 2). Constructions are combined when a unit, potentially itself composed of constructions, fills a slot in another construction. Viewing language as a system of constructions rather than words and rules may fundamentally change how the successes and failures of models are construed, and new goals and questions come into focus. The perspective on language motivates a move away from certain traditional evaluation practices. Specifically, we recommend the following.
When possible, use a variety of naturalistic sentences rather than sentences generated by a template that presupposes grammatical rules with interchangeable vocabulary items. The idea that sentences can be constructed by subbing random lexical items into templates often misses lexical subtleties that are an important part of natural language. Instead, ecologically valid stimuli can be collected or adapted from natural corpora and normed for naturalness and plausibility. Since human judgments are highly context-dependent, benchmark tasks should also vary contexts systematically (see, e.g., Ross et al., 2024; Parrish et al., 2021).
Collect human assessments that allow for gradient context-sensitive interpretations that appeal to learned constructions. Evaluating LM competence on individual constructions requires assessing both acceptability judgments and interpretations from humans to make fair comparisons.
We also need to be sensitive to the implications people and LMs draw from instructions and the testing context. Do not give instructions to human-evaluators (or models) in ways that make the results a foregone conclusion. If people are instructed to interpret ‘red X’ as ‘X that is red for any X,’ they are capable of doing so; this may reflect the instructions, not their natural intuitions. In natural contexts, people understand that red grapefruits are closer to pink, red hair is more orange, a red book may be about communism, and crossing a red line may have consequences.
The items included in testing also influence interpretations by people and models, by providing context. For instance, if certain pairs of items are mismatched (e.g., “The cup is green.” “The cup is blue.”) while others are matched, people can infer which ones are intended to be contradictory. LMs, like people, are now capable of generalizing by rule when tasked to do so. For instance, Lampinen et al. (2025) found Gemini 1.5 Flash Team et al. (2024) avoided the reversal curse, achieving 100% accuracy, when the Berglund et al. (2023) dataset was provided to the LM as context.
A variety of items, participants, and contexts ought to be valued as much as a variety of models. Because linguistic meaning is deeply tied to local context, even seemingly similar sentences can have very different interpretations in ways that depend on context. Therefore, experiments that vary context are desirable.
The most interesting questions may no longer be whether LMs are skilled at producing and responding to natural languages, but how they achieve such remarkable skills. As is familiar from the lexicon, constructions comprise an interrelated network. We can now study relationships among constructions mechanistically. For instance, Misra and Mahowald (2024) have demonstrated that even when all instances of a rare non-compositional construction are ablated from training data, non-trivial learning of the construction remains, enabled by the presence of related constructions in training. Moreover, nearly every productive construction coexists with at least a few formulaic instances, and LMs offer ways to test various theoretical perspectives on the nature of those relationships. These and other newer ways of probing LMs are possible, and our toolkit will only grow.
We can investigate whether LMs distinguish subtle meaningful differences between constructions Tseng et al. (2022); Li et al. (2022); Weissweiler et al. (2022); Tayyar Madabushi et al. (2020); Devlin et al. (2019); Zhou et al. (2024). How they do so or when and why they fall short is another important direction.
8 Conclusion
Generalization is a key component of human language—and a big part of why LMs are successful at processing language. But we have argued that evaluations of the linguistic abilities of LMs are too often based on an assumption that generalization requires algebraic rules operating on words. Natural languages are not Lego sets. Instead, language involves flexible combinations of rich and varied constructions of differing sizes, complexities, and degrees of abstraction, which differ from algebraic rules in many ways. By designing new evaluations that accurately reflect the complexities of language, we can avoid under- or overestimating language models. The extent to which LMs produce and interpret combinations of constructions has only begun to be explored. We believe future progress lies not in asking whether LMs obey abstract rules, but in probing what kinds of constructions they learn, how they relate them, and how those structures guide novel interpretation and production. In doing so, we may better capture what it truly means to comprehend and use language.
Limitations
While we have aimed to discuss benchmarks and evaluations in ways that reflect the historical trajectory as well as the present-day landscape, evaluations of LMs are continually developing. We feel the dominant paradigms have and continue to be based on data generated by rules and evaluated without regard for context effects, gradience, or semantic nuance, but we are keenly aware that we have likely overlooked metrics that go beyond rule-based evaluations (e.g., Parrish et al., 2021).
We recognize the growing work in multilingual evaluations, which are inherently valuable Mueller et al. (2020); Jumelet et al. (2025); Kryvosheieva and Levy (2025). The current perspective applies to all natural languages, but comparative work is not the focus of the current perspective, and we use English examples for the sake of easy comprehension and brevity.
We have taken pains to explain the appeal of constructions over rules (Table 1 and §5), but since constructions may be less familiar, we only hope that our brief perspective may whet readers’ appetite to learn more.
Acknowledgments
We thank Kanishka Misra and Will Merrill for helpful discussions and feedback. We are grateful to audiences at the NSF-sponsored New Horizons in Language Science workshop and the Analytical approaches to understanding neural networks summmer school sponsored by Simon’s Foundation for helpful feedback. Leonie Weissweiler was supported by a postdoctoral fellowship of the German Academic Exchange Service (DAAD).